# A Novel Hierarchical Co-Optimization Framework for Coordinated Task Scheduling and Power Dispatch in Computing Power Networks

### v1

@@ Abstract
The proliferation of large-scale artificial intelligence and data-intensive applications has spurred the development of Computing Power Networks (CPNs), which promise to deliver ubiquitous and on-demand computational resources. However, the immense energy consumption of these networks poses a significant sustainability challenge. Simultaneously, power grids are grappling with the instability introduced by the high penetration of intermittent renewable energy sources (RES). This paper addresses these dual challenges through a novel Two-Stage Co-Optimization (TSCO) framework that synergistically manages power system dispatch and CPN task scheduling to achieve low-carbon operations. The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment (SUC) stage and a real-time operational stage. The former is solved using Benders decomposition for computational tractability, while in the latter, economic dispatch of generation assets is coupled with an adaptive CPN task scheduling managed by a Deep Reinforcement Learning (DRL) agent. This agent makes intelligent, carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity. Through extensive simulations on an IEEE 30-bus system integrated with a CPN, the TSCO framework is shown to significantly outperform baseline approaches. Results demonstrate that the proposed framework reduces total carbon emissions and operational costs, while simultaneously decreasing RES curtailment by more than $60\%$ and maintaining stringent Quality of Service (QoS) for computational tasks.
Index Terms:
Computing Power Network (CPN), renewable energy, data center, carbon-aware scheduling, deep reinforcement learning.
@@ I-B Research Motivation
The central challenge lies in a fundamental misalignment between the operational dynamics of CPNs and RES-dominated power grids. CPNs are designed to host a wide array of computational tasks, from latency-sensitive services like cloud virtual reality to compute-intensive, batch-processing jobs like large model training [8], [9]. These tasks demand a highly reliable and stable power supply to ensure Quality of Service (QoS). However, RES are inherently variable, non-dispatchable, and dependent on weather conditions. This intermittency introduces significant volatility into the power grid, leading to challenges in maintaining frequency and voltage stability, and often results in the wasteful curtailment of clean energy when generation exceeds demand [10], [11].
Simply connecting a CPN to a grid with high RES penetration without intelligent coordination creates a direct conflict. During periods of low RES output, the CPN’s demand would force the grid to rely on expensive and carbon-intensive fossil-fuel peaker plants to maintain balance. Conversely, during periods of high RES output, the grid may be forced to curtail wind or solar generation to prevent overload, even as CPNs continue to draw power from a mixed-carbon source. This decoupled operation leads to a suboptimal outcome: either unreliable computation or high operational costs and carbon emissions. The problem is not merely about sourcing green energy, but about managing the flexibility of CPN workloads as a grid-stabilizing asset. The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [12]. This reframes the problem from ”how to power the CPN cleanly” to ”how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”. This symbiotic relationship is the core motivation for the co-optimization framework proposed in this paper.
@@ I-C Contributions of the Paper
  • A Comprehensive Integrated System Model: A detailed model is developed that captures the intricate interplay between a heterogeneous CPN and a modern power grid. The model incorporates conventional thermal generators, stochastic RES, and battery energy storage systems (BESS), while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
@@ II Related Works
This section provides a structured review of the literature across four key domains that intersect in this work, Computing Power Network (CPN) architectures, integrated energy system management, carbon-aware workload scheduling, and optimization under uncertainty in power systems. This analysis serves to contextualize our contribution and highlight the research gap that our proposed framework aims to fill.
@@ II-A Computing Power Network Architectures and Scheduling
The concept of the CPN has evolved from earlier paradigms like cloud, fog, and edge computing, with the primary goal of creating a unified network for ubiquitous computing resources. Early research focused on defining the architecture and core features of CPNs, such as intent-driven operation, closed-loop autonomy, and elastic scheduling. Architectures have been proposed with both centralized control planes, which possess a global view for unified scheduling, and distributed schemes, where decisions are made locally by network nodes.
A significant body of research in CPNs has concentrated on task scheduling [13]. The primary objectives have traditionally been to optimize QoS metrics. For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques. Other works have explored task offloading in terminal-side CPNs or the joint selection of routing paths and computing nodes [14]. While some research considers energy consumption as a constraint or a secondary objective, the explicit, primary optimization of carbon emissions based on the real-time state of the power grid remains largely unexplored. Existing CPN scheduling literature typically treats the power grid as an external, static entity, overlooking the potential for dynamic, symbiotic interaction.
@@ II-B Integrated Energy System Management
In the power systems domain, the concept of coordinating generation, transmission, and consumption has been studied extensively under the ”Generation-Grid-Load-Storage” integrated operation model [11]. This paradigm seeks to improve system safety, economy, and reliability through the coordinated interaction of all system components. Research in this area includes the development of multi-timescale optimal dispatching strategies [15], economic dispatch models that incorporate DR, and optimal power flow (OPF) formulations that aim to minimize generation costs while respecting network constraints [16].
The integration of flexible loads and DR has been identified as a key enabler for grids with high RES penetration. Studies have explored how to coordinate data centers as flexible loads with a load aggregator to minimize electricity costs and absorb grid volatility [17]. However, these studies often rely on simplified models of the flexible load, such as an abstract ability to shift power consumption in time, without capturing the complex internal constraints, dependencies, and heterogeneous resource requirements of a CPN workload. They treat the CPN as a ”black box” load, missing the opportunity to optimize its internal operations in concert with the grid.
@@ II-C Carbon-Aware Workload Scheduling
The field of green computing has produced a significant body of work on carbon-aware scheduling for geo-distributed data centers. A common strategy is spatio-temporal scheduling [18], which involves shifting computational workloads in time (”temporal flexibility”) or space (”geographical flexibility”) to data centers with lower electricity prices or cleaner energy mixes. These methods often leverage real-time carbon intensity signals from services like WattTime [^2] https://watttime.org/data-science/data-signals/ to guide scheduling decisions.
To manage carbon emissions over the long term, some works have proposed online algorithms based on Lyapunov optimization [19]. This technique transforms a long-term average constraint, e.g., a carbon budget, into a series of real-time optimization subproblems by maintaining a ”virtual queue” that tracks the deviation from the budget. The scheduler is then penalized for actions that increase this queue length [20]. While powerful, this line of research suffers from a critical limitation. It almost universally treats the power grid as an exogenous system. The price and carbon intensity signals are assumed to be external inputs that are unaffected by the scheduling decisions. This assumption breaks down at scale, as the collective actions of large CPNs can and will influence grid operations, market prices, and the generation mix, thereby altering the very signals they are responding to.
@@ III System Architecture and Problem Formulation
This section presents the mathematical models for the CPN and the integrated power system. These models form the basis of our co-optimization framework, as shown in Fig. 1.
@@
III-A Computing Power Network (CPN) Model
@@ III-A1 CPN Node Model
Each CPN node $n\in\mathcal{N}$ is characterized by its available computational resources, such as CPUs, GPUs, and TPUs. For simplicity in this formulation, we aggregate these into a single measure of computing capacity, $C_{n}^{comp}$, expressed in Floating Point Operations Per Second (FLOPS). The power consumption of a CPN node is modeled as a linear function of its utilization [21], varying between an idle power $P_{n}^{idle}$ and a peak power $P_{n}^{peak}$ at full utilization. The total power consumed by node $n$ at time $t$ is a decision variable, $P_{n,t}^{CPN}$, determined by the scheduled tasks.
@@ III-D Joint Optimization Problem Formulation
    $\sum_{n\in\mathcal{N}}\sum_{t=A_{k}}^{D_{k}}x_{k,\tau,n,t}=1\quad\forall k,\tau\in\mathcal{T}_{k}.$ (7)
@@
    $\sum_{k\in\mathcal{J}}\sum_{\tau\in\mathcal{T}_{k}}r_{\tau}x_{k,\tau,n,t}\leq C_{n}^{comp}\quad\forall n,t.$ (8)
  • Coupling Constraint: This is the critical link. The power consumed by a CPN node $n$ (co-located at bus $i$) is determined by the tasks scheduled on it.
    $P_{i,t,\omega}^{CPN}=P_{n}^{idle}+(P_{n}^{peak}-P_{n}^{idle})\frac{\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}}{C_{n}^{comp}}.$ (9)
    This equation makes the CPN load term $P_{i,t,\omega}^{CPN}$ in the power balance equation an endogenous variable dependent on the scheduling decisions $x_{k,\tau,n,t}$.
@@ IV-A Hierarchical Stochastic Optimization Structure
      Real-Time Economic Dispatch (ED): Solves for the optimal power output of the committed generators and the charge/discharge schedule of BESS to meet the actual load at minimum cost, while respecting all grid constraints.

### v2

@@ Abstract
The proliferation of large-scale artificial intelligence and data-intensive applications has spurred the development of Computing Power Networks (CPNs), which deliver ubiquitous and on-demand computational resources. However, the immense energy consumption poses a significant sustainability challenge. Simultaneously, power grids are grappling with the instability by the high penetration of intermittent renewable energy sources (RES). This paper addresses these dual challenges through a novel Two-Stage Co-Optimization (TSCO) framework that synergistically manages power system dispatch and CPN task scheduling to achieve low-carbon operations. The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment stage and a real-time operational stage. The former is solved using Benders decomposition for computational tractability, while in the latter, economic dispatch of generation assets is coupled with an adaptive CPN task scheduling managed by a deep reinforcement learning agent. It makes carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity. Through extensive simulations on an IEEE 30-bus system integrated with a CPN, the TSCO framework is shown to significantly outperform baseline approaches. It can reduce carbon emissions and operational costs, while simultaneously decreasing RES curtailment by more than $60\%$ and maintaining a stringent quality of service for computational tasks.
@@ I-B Research Motivation
The central challenge lies in a fundamental misalignment between the operational dynamics of CPNs and RES-dominated power grids. CPNs are designed to host a wide array of computational tasks, from latency-sensitive services like cloud virtual reality to compute-intensive, batch-processing jobs like large model training [8], [9]. These tasks demand a highly reliable and stable power supply to ensure Quality of Service (QoS). However, RES are inherently variable, non-dispatchable, and dependent on weather conditions. This intermittency introduces significant volatility into the power grid, leading to challenges in maintaining frequency and voltage stability, and often results in the wasteful curtailment of clean energy when generation exceeds demand [10].
Simply connecting a CPN to a grid with high RES penetration without intelligent coordination creates a direct conflict. During periods of low RES output, the CPN’s demand would force the grid to rely on expensive and carbon-intensive fossil-fuel peaker plants to maintain balance. Conversely, during periods of high RES output, the grid may be forced to curtail wind or solar generation to prevent overload, even as CPNs continue to draw power from a mixed-carbon source. This decoupled operation leads to a suboptimal outcome: either unreliable computation or high operational costs and carbon emissions. The problem is not merely about sourcing green energy, but about managing the flexibility of CPN workloads as a grid-stabilizing asset. The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [11]. This reframes the problem from “how to power the CPN cleanly” to “how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”. This symbiotic relationship is the core motivation for the co-optimization framework proposed in this paper.
@@ I-C Contributions of the Paper
  • A Comprehensive Integrated System Model: A detailed model is developed that captures the intricate interplay between a heterogeneous CPN and a modern power grid. It incorporates conventional thermal generators, stochastic RES, and battery energy storage systems (BESS), while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
@@ II Related Works
This section provides a structured review of the literature across three key domains that intersect in this work, including Computing Power Network (CPN) architectures, integrated energy system management, and carbon-aware workload scheduling. This analysis serves to contextualize our contribution and highlight the research gap that our proposed framework aims to fill.
@@ II-A Computing Power Network Architectures and Scheduling
The CPN concept has evolved from earlier paradigms, such as cloud, fog, and edge computing, with the primary goal of creating a unified network for ubiquitous computing resources. Early research focused on defining the architecture and core features of CPNs, such as intent-driven operation, closed-loop autonomy, and elastic scheduling. Architectures have been proposed with both centralized control planes, which possess a global view for unified scheduling, and distributed schemes, where decisions are made locally by network nodes.
A significant body of research in CPNs has concentrated on task scheduling [12]. The primary objectives have traditionally been to optimize QoS metrics. For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques. Other works have explored task offloading in terminal-side CPNs [13] or the joint selection of routing paths and computing nodes [14]. While some research considers energy consumption as a constraint or a secondary objective, the explicit, primary optimization of carbon emissions based on the real-time state of the power grid remains largely unexplored. Existing CPN scheduling literature typically treats the power grid as an external, static entity, overlooking the potential for dynamic, symbiotic interaction.
@@ II-B Integrated Energy System Management
In the power systems domain, the concept of coordinating generation, transmission, and consumption has been studied extensively under the “Generation-Grid-Load-Storage” integrated operation model [15]. This paradigm seeks to improve system safety, economy, and reliability through the coordinated interaction of all components. Research in this area includes the development of multi-timescale optimal dispatching strategies [16], economic dispatch models that incorporate DR, and Optimal Power Flow (OPF) formulations that aim to minimize generation costs while respecting network constraints [17].
The integration of flexible loads and DR has been identified as a key enabler for grids with high RES penetration. Studies have explored how to coordinate data centers as flexible loads with a load aggregator to minimize electricity costs and absorb grid volatility [18]. However, these studies often rely on simplified models of the flexible load, such as an abstract ability to shift power consumption in time, without capturing the complex internal constraints, dependencies, and heterogeneous resource requirements of a CPN workload. They treat the CPN as a “closed box” load, missing the opportunity to optimize its internal operations in concert with the grid [19].
@@ II-C Carbon-Aware Workload Scheduling
The green computing field has produced a substantial body of work on carbon-aware scheduling for geographically distributed data centers. A common strategy is spatio-temporal scheduling [20], which involves shifting computational workloads in time or space to data centers with lower electricity prices or cleaner energy mixes. These methods often leverage real-time carbon intensity signals from services like WattTime [^2] https://watttime.org/data-science/data-signals/ to guide scheduling decisions.
To manage carbon emissions over the long term, some works have proposed online algorithms based on Lyapunov optimization [21]. This technique transforms a long-term average constraint, e.g., a carbon budget, into a series of real-time optimization subproblems by maintaining a “virtual queue” that tracks the deviation from the budget. The scheduler is then penalized for actions that increase this queue length [22]. While powerful, this line of research suffers from a critical limitation. It almost universally treats the power grid as an exogenous system [23]. The price and carbon intensity signals are assumed to be external inputs that are unaffected by the scheduling decisions. This assumption breaks down at scale, as the collective actions of large CPNs can and will influence grid operations, market prices, and the generation mix, thereby altering the very signals they are responding to.
@@ III System Architecture and Problem Formulation
This section presents the mathematical models for the CPN and the integrated power system. These models form the basis of our co-optimization framework, as shown in Fig. 1. For clarity, Table I summarizes the key notations used throughout this paper.
@@
TABLE I: Summary of Key Notations
Symbol Description
$\mathcal{N}$ Set of CPN nodes (data centers)
$\mathcal{G}_{C}$ Set of conventional thermal generators
$\mathcal{G}_{R}$ Set of renewable energy (RES) generators (wind/solar)
$\mathcal{J}$ Set of computational jobs in CPN
$\mathcal{T}_{k}$ Set of subtasks of job $k$
$\mathcal{I}$ Set of buses in the IEEE 30-bus power system
$C_{n,h}^{comp}$ Computing capacity of hardware $h$ at CPN node $n$ (in FLOPS)
$P_{n,t}^{CPN}$ Total power consumption of CPN node $n$ at time $t$
$P_{g,t}$ Power output of conventional generator $g$ at time $t$ (in MWh)
$P_{g,t,\omega}^{R,avail}$ Available power output of RES generator $g$ at time $t$ under scenario $\omega$
$x_{k,\tau,n,h,t}$ Binary variable: 1 if subtask $\tau$ of job $k$ is assigned to hardware $h$ at node $n$ and time $t$
$u_{g,t}$ Commitment status of conventional generator $g$ at time $t$ (1=on, 0=off)
$SOC_{b,t,\omega}$ State-of-charge of battery energy storage system (BESS) $b$ at time $t$ under scenario $\omega$
$\epsilon_{g}$ Carbon emission factor of generator $g$ (tons of $\text{CO}_{2}$ per MWh)
$E_{budget}$ Long-term carbon emission budget for the planning horizon
$LMP_{i,t}$ Locational marginal price at bus $i$ and time $t$ (in \$/MWh)
$MCI_{i,t}$ Marginal carbon intensity at bus $i$ and time $t$
$\Omega$ Set of uncertainty scenarios
III-A Computing Power Network Model
@@ III-A1 CPN Node Model
Each CPN node $n\in N$ consists of heterogeneous hardware resources, denoted by the set $H=\{CPU,GPU,TPU\}$. For each hardware type $h\in H$, we define its key parameters: $C_{n,h}^{comp}$ is computing capacity in FLOPS); $P_{n,h}^{idle}$ is idle power consumption; $P_{n,h}^{peak}$ denotes peak power consumption at full utilization; $\alpha_{n,h},\beta_{n,h}$ are nonlinear coefficients calibrated from real-world hardware traces.
The power consumption of hardware $h$ at node $n$ follows a quadratic nonlinear model, capturing the superlinear power growth of accelerators (e.g., GPU/TPU) under high utilization:
$P_{n,h,t}^{comp}=P_{n,h}^{idle}+(\alpha_{n,h}\cdot u_{n,h,t}^{2}+\beta_{n,h}\cdot u_{n,h,t})\cdot(P_{n,h}^{peak}-P_{n,h}^{idle}),$ (1)
where $u_{n,h,t}$ is the utilization of hardwar $h$ at node $n$ and time $t$, calculated as:
$u_{n,h,t}=\frac{\sum_{k\in\mathcal{J}}\sum_{\tau\in\mathcal{T}_{k}}r_{\tau,h}\cdot x_{k,\tau,n,h,t}}{C_{n,h}^{comp}},$ (2)
where $r_{\tau,h}$ denotes the resource requirement of subtask $\tau$ for hardwarehh, set to 0 if $\tau$ does not require $h$. $x_{k,\tau,n,h,t}$ is a binary variable. Specifically, $1$ if subtask $\tau$ of job $k$ is assigned to hardware $h$ at node $n$ and time $t$, 0 otherwise. The total power consumption of nodenn and time $t$ is the sum of power from all hardware types:
$P_{n,t}^{CPN}=\sum_{h\in H}P_{n,h,t}^{comp}.$ (3)
@@ III-A2 Task Model
Each subtask $\tau\in T_{k}$ is further defined by: (1) hardware type constraint $h_{\tau}\subseteq H$, the set of hardware types capable of executing $\tau$, e.g., large-model training subtasks require $h_{\tau}=\{GPU,TPU\}$; (2) data volume $d_{\tau}$, in GB, for communication constraint calculation; (3) total computational workload $w_{\tau}$ in FLOPs, and resource requirement $r_{\tau,h}$ is the number of processing units for hardware $h\in h_{\tau}$. The execution time of subtask $\tau$ on hardware $h$ of node $n$ is:
$t_{\tau,n,h}=\begin{cases}\frac{w_{\tau}}{C_{n,h}^{\mathrm{comp}}},&\text{if }h\in h_{\tau}\\ \infty,&\text{otherwise}\end{cases}\quad\forall n,h\in h_{\tau},\tau\in\mathcal{T}_{k}.$ (4)
III-A3 Communication Link Model
The communication network between CPN nodes is modeled as a directed graph $\mathcal{G}_{comm}=(\mathcal{N},\mathcal{L})$, where $\mathcal{L}$ denotes the set of bidirectional communication links between nodes. For each link from the source node to the destination node $(n_{src},n_{dest})\in\mathcal{L}$, two key parameters are defined based on real-world network traces:
  • Bandwidth: $B_{n_{src},n_{dest}}$ (GB/s) is the maximum data transmission rate of the link, fixed for dedicated CPN networks.
  • Latency parameters: It includes $lat_{n_{src},n_{dest}}^{base}$ (s, base latency without data transmission) and $\gamma_{n_{src},n_{dest}}$ (s/GB, dynamic latency coefficient related to data volume). Then, the total latency for transmitting a subtask $\tau$ with data volume $d_{\tau}$ from $n_{src}$ to $n_{dest}$ is:
    $lat_{n_{src},n_{dest}}(\tau)=lat_{n_{src},n_{dest}}^{base}+\gamma_{n_{src},n_{dest}}\cdot d_{\tau}.$ (5)
    The data migration time (i.e., the duration of data transmission) is determined by link bandwidth:
    $t_{migrate,\tau,n_{src},n_{dest}}=\frac{d_{\tau}}{B_{n_{src},n_{dest}}}.$ (6)
@@ III-D Joint Optimization Problem Formulation
    $\sum_{n\in\mathcal{N}}\sum_{h\in\mathcal{h_{\tau}}}\sum_{t=A_{k}}^{D_{k}}x_{k,\tau,n,h,t}=1\quad\forall k,\tau\in\mathcal{T}_{k}.$ (13)
@@
    $\sum_{k\in\mathcal{J}}\sum_{\tau\in\mathcal{T}_{k}}r_{\tau,h}x_{k,\tau,n,h,t}\leq C_{n,h}^{comp}\quad\forall n,h\in h_{\tau},t.$ (14)
    Migration Transmission Time Constraint: For any subtask $\tau$ migrated from node $n_{src}$ to $n_{dest}$ (i.e., $x_{k,\tau,n_{src},h,t_{1}}=0$ and $x_{k,\tau,n_{dest},h,t_{2}}=1$ with $t_{2}>t_{1}$), the start time of $\tau$ at $n_{dest}$ must account for data migration time:
    $start_{k,\tau,n_{dest},h,t_{2}}\geq$ $finish_{k,\tau_{pre},n_{src},h^{\prime},t_{1}}+$ (15)
    $t_{migrate,\tau,n_{src},n_{dest}},$
    $\forall(\tau_{pre},\tau)\in E_{k},h\in h_{\tau},h^{\prime}\in h_{\tau_{pre}},$ (16)
    where $finish_{k,\tau_{pre},n_{src},h^{\prime},t_{1}}$ is the completion time of predecessor subtask $\tau_{pre}$ at $n_{src}$, and $t_{migrate,\tau,n_{src},n_{dest}}$ is calculated via Eq. (6).
    Link Bandwidth Constraint: For each communication link $(n_{src},n_{dest})\in\mathcal{L}$, the total data transmission rate of all concurrent migration tasks must not exceed the link bandwidth:
    $\sum_{\begin{subarray}{c}migrate(\tau,n_{src}\to n_{dest})\end{subarray}}\frac{d_{\tau}}{t_{migrate,\tau,n_{src},n_{dest}}}\leq B_{n_{src},n_{dest}},$ (17)
    $\forall t,(n_{src},n_{dest})\in\mathcal{L},$ (18)
    were $migrate(\tau,n_{src}\to n_{dest})$ denotes that subtask $\tau$ is migrated from $n_{src}$ to $n_{dest}$, and $\tau^{\prime}$ is the time interval of data transmission.
  • Coupling Constraint: The total power consumption of CPN node $n$ co-located at bus $i$ is determined by the nonlinear power model of heterogeneous hardware and scheduling decisions:
    $P_{i,n,t}^{CPN}=\sum_{h\in h_{\tau}}[P_{n,h}^{idle}+$ $(\alpha_{n,h}\cdot u_{n,h,t}^{2}+\beta_{n,h}\cdot u_{n,h,t})\cdot$ (19)
    $(P_{n,h}^{peak}-P_{n,h}^{idle})].$
    The CPN power consumption is a nonlinear function, while the real-time economic dispatch (ED) of the power grid is a linear programming problem. The direct coupling of these two will lead to a significant increase in the computational complexity of the optimization problem. To balance accuracy and tractability, we can use piecewise linearization to handle the nonlinear terms [30].
@@ IV-A Hierarchical Stochastic Optimization Structure
      Real-Time ED: Solves for the optimal power output of the committed generators and the charge/discharge schedule of BESS to meet the actual load at minimum cost, while respecting all grid constraints.
@@
The interface between the two stages is designed to ensure tight coupling while maintaining computational tractability. Specifically, Stage 1 outputs three types of binding constraints to Stage 2: (1) the commitment status $u_{g,t}^{*}$ of conventional generators, which restricts the set of dispatchable units in real-time ED; (2) the energy reservation bounds $(SOC_{b,t}^{reserve,min},SOC_{b,t}^{reserve,max})$ for BESS, ensuring sufficient flexibility to accommodate real-time RES volatility; (3) the per-node computational resource reservation $C_{n,t}^{comp,reserve}$ for CPN, which limits the maximum resource utilization of each node to avoid overloading beyond the day-ahead planning envelope. In return, Stage 2 feeds back key real-time operational data to Stage 1 for iterative improvement. It includes the deviation between actual and forecasted RES generation, the CPN task completion rate and deadline adherence, and the cumulative carbon emissions up to the current time. This feedback is used to refine the scenario set $\Omega$ for the next day’s SUC problem, e.g., adjusting the probability of high RES curtailment scenarios based on real-time absorption capacity. The interface operates on a daily rolling horizon. The day-ahead plan is updated once every 24 hours, while the real-time stage fetches the latest day-ahead constraints every 5-15 minutes and reports cumulative operational data at the end of each day.